Operation Summary for August 18, 2010 - August 20, 2010 |
Wednesday August 18 The day shift began with startup activities in the Tevatron (TeV), Recycler, Antiproton Source (Pbar), Linac, Booster, Main Injector (MI), NuMI, MiniBooNE, Switchyard (SY), and the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments. During the midnight shift, operators took the TeV separators out of condition mode. The TeV sector A3 air compressor tripped off; the Duty Electrician reported a bad contactor. Operators prepared MI, Booster, NuMI, MINOS, 8GeV line, Pbar rings, and the TeV enclosures for access. During the day shift, the TeV checkout began at 10 AM. Operations reported that the MI had to be searched and secured several times due to numerous dropped sectors. Booster accesses concluded at 2:25 PM. During the evening shift, Operators searched and secured enclosures B, C, D, E and MI-65. Operators turned on the MI-65 power supplies for ramp testing. The MI personnel completed their access by 6:20 PM. Operations established circulating beam to the MI dump by 9:20 PM. Thursday August 19 The midnight shift began startup preparations. During the midnight shift, Operations reported that MI-8 line studies occurred for the first half of the shift. Operators switched on the SY power supplies HP3DS and QP3. Interlock technicians began testing the SY safety system from 6 AM to 7:30 AM. Operations switched off the MI-65 power supplies and dropped the enclosure interlocks for scheduled day shift activities. Operations also dropped the CDF interlocks. During the day shift, start-up activities continued. Operations conducted beam commissioning for MI, Pbar, and Recycler. Safety personnel performed a safety walk-through of TeV. By the end of the shift, accesses were underway in MI for a hot magnet investigation and F-Sector for a ground fault on a P150 line magnet string. During the evening shift, MI personnel reported finding a closed LCW valve. Operations established beam to the MI abort and Pbar. Operations turned on the TeV power supplies while the D4 cryo system continued to cool down. D4 was cold by 11:30 PM and TeV checkout began. Operations had current on the TeV bus by the end of the shift. NuMI target scans were put on hold due to power supply problems with a Lambertson (LAM60) and LCW temperature issues at MI-62. Friday August 20 The midnight shift began startup preparations. During the midnight shift, the TeV quenched on the first HEP ramp. Operations recovered the TeV and had current re-established to the bus by 2:45 AM. Operations reported problems with the Booster 400 MeV area Debuncher LCW temperature control system. TeV studies began at 6:30 AM. The Plans for Friday and the Weekend The plans for today are to continue startup activities and try a TeV shot on Saturday. |
|
More Information Comments and Suggestions |
| last modified 08/20/2010 email Fermilab |
| Security, Privacy, Legal | |